# Disentangling Reasoning and Knowledge in Medical Large Language Models

### v1

@@ Appendix B Additional Data Curation Details
MedReason-8B 61.259.9 - ${}_{\text{\tiny 59.9 - 62.5}}$ 44.742.9 - ${}_{\text{\tiny 42.9 - 46.4}}$ 61.158.3 - ${}_{\text{\tiny 58.3 - 63.7}}$ 32.731.0 - ${}_{\text{\tiny 31.0 - 34.4}}$ 61.259.9 - ${}_{\text{\tiny 59.9 - 62.3}}$ 39.138.0 - ${}_{\text{\tiny 38.0 - 40.2}}$
m1-7B-23k 67.265.9 - ${}_{\text{\tiny 65.9 - 68.5}}$ 49.647.8 - ${}_{\text{\tiny 47.8 - 51.4}}$ 68.566.0 - ${}_{\text{\tiny 66.0 - 70.9}}$ 38.136.3 - ${}_{\text{\tiny 36.3 - 40.0}}$ 67.566.3 - ${}_{\text{\tiny 66.3 - 68.5}}$ 44.242.9 - ${}_{\text{\tiny 42.9 - 45.4}}$
HuatuoGPT-o1-8B (SFT) 62.761.4 - ${}_{\text{\tiny 61.4 - 64.0}}$ 46.945.3 - ${}_{\text{\tiny 45.3 - 48.4}}$ 63.060.6 - ${}_{\text{\tiny 60.6 - 65.6}}$ 33.331.6 - ${}_{\text{\tiny 31.6 - 35.1}}$ 62.761.6 - ${}_{\text{\tiny 61.6 - 64.0}}$ 40.539.2 - ${}_{\text{\tiny 39.2 - 41.7}}$
HuatuoGPT-o1-8B (RL) 63.261.9 - ${}_{\text{\tiny 61.9 - 64.5}}$ 46.745.0 - ${}_{\text{\tiny 45.0 - 48.4}}$ 69.666.9 - ${}_{\text{\tiny 66.9 - 72.1}}$ 34.332.6 - ${}_{\text{\tiny 32.6 - 35.8}}$ 64.563.3 - ${}_{\text{\tiny 63.3 - 65.7}}$ 40.839.6 - ${}_{\text{\tiny 39.6 - 42.1}}$
BioMed-R1-8B (SFT) 67.666.2 - ${}_{\text{\tiny 66.2 - 68.8}}$ 49.247.5 - ${}_{\text{\tiny 47.5 - 50.9}}$ 70.868.2 - ${}_{\text{\tiny 68.2 - 73.4}}$ 40.038.2 - ${}_{\text{\tiny 38.2 - 41.7}}$ 68.267.0 - ${}_{\text{\tiny 67.0 - 69.2}}$ 44.943.6 - ${}_{\text{\tiny 43.6 - 46.1}}$
BioMed-R1-8B (RL) 70.068.8 - ${}_{\text{\tiny 68.8 - 71.2}}$ 53.351.5 - ${}_{\text{\tiny 51.5 - 55.0}}$ 72.069.5 - ${}_{\text{\tiny 69.5 - 74.4}}$ 39.337.6 - ${}_{\text{\tiny 37.6 - 41.1}}$ 70.469.3 - ${}_{\text{\tiny 69.3 - 71.5}}$ 46.745.4 - ${}_{\text{\tiny 45.4 - 47.9}}$
Table S7: In-distribution (ID) and out-of-distribution (OOD) accuracy for biomedical reasoning models across knowledge-heavy, reasoning-heavy, and overall subsets. Results are shown with 95% confidence intervals. BioMed-R1-8B (RL) achieves the strongest overall and OOD performance, suggesting reinforcement learning improves generalization, particularly on knowledge-based questions. In contrast, gains in reasoning accuracy are mostly concentrated in-distribution.

### v2

@@ Appendix B Additional Training Details
Llama w/ m23k SFT 67.265.9 - ${}_{\text{\tiny 65.9 - 68.5}}$ 49.647.9 - ${}_{\text{\tiny 47.9 - 51.2}}$ 68.565.8 - ${}_{\text{\tiny 65.8 - 71.2}}$ 38.136.4 - ${}_{\text{\tiny 36.4 - 39.8}}$ 67.566.3 - ${}_{\text{\tiny 66.3 - 68.7}}$ 44.242.9 - ${}_{\text{\tiny 42.9 - 45.3}}$
+ Random Hard SFT 66.865.5 - ${}_{\text{\tiny 65.5 - 68.1}}$ 49.047.3 - ${}_{\text{\tiny 47.3 - 50.7}}$ 70.367.6 - ${}_{\text{\tiny 67.6 - 72.9}}$ 38.536.7 - ${}_{\text{\tiny 36.7 - 40.3}}$ 67.466.2 - ${}_{\text{\tiny 66.2 - 68.6}}$ 44.042.8 - ${}_{\text{\tiny 42.8 - 45.3}}$
+ Knowledge Hard SFT 67.265.9 - ${}_{\text{\tiny 65.9 - 68.4}}$ 48.847.2 - ${}_{\text{\tiny 47.2 - 50.4}}$ 69.667.1 - ${}_{\text{\tiny 67.1 - 72.1}}$ 38.436.8 - ${}_{\text{\tiny 36.8 - 40.1}}$ 67.666.5 - ${}_{\text{\tiny 66.5 - 68.7}}$ 43.942.7 - ${}_{\text{\tiny 42.7 - 45.2}}$
+ Reason Hard SFT 67.666.3 - ${}_{\text{\tiny 66.3 - 68.9}}$ 49.247.6 - ${}_{\text{\tiny 47.6 - 50.9}}$ 70.868.1 - ${}_{\text{\tiny 68.1 - 73.2}}$ 40.038.1 - ${}_{\text{\tiny 38.1 - 41.6}}$ 68.267.1 - ${}_{\text{\tiny 67.1 - 69.3}}$ 44.943.6 - ${}_{\text{\tiny 43.6 - 46.1}}$
+ Random Hard RL 69.267.8 - ${}_{\text{\tiny 67.8 - 70.4}}$ 52.550.8 - ${}_{\text{\tiny 50.8 - 54.2}}$ 71.769.1 - ${}_{\text{\tiny 69.1 - 74.2}}$ 39.437.6 - ${}_{\text{\tiny 37.6 - 41.1}}$ 69.668.5 - ${}_{\text{\tiny 68.5 - 70.8}}$ 46.345.0 - ${}_{\text{\tiny 45.0 - 47.6}}$
+ Knowledge Hard RL 69.668.4 - ${}_{\text{\tiny 68.4 - 70.9}}$ 52.150.5 - ${}_{\text{\tiny 50.5 - 53.9}}$ 72.670.1 - ${}_{\text{\tiny 70.1 - 74.9}}$ 39.838.0 - ${}_{\text{\tiny 38.0 - 41.4}}$ 70.269.1 - ${}_{\text{\tiny 69.1 - 71.3}}$ 46.345.0 - ${}_{\text{\tiny 45.0 - 47.6}}$
+ Reason Hard RL 70.068.7 - ${}_{\text{\tiny 68.7 - 71.2}}$ 53.351.5 - ${}_{\text{\tiny 51.5 - 55.0}}$ 72.069.4 - ${}_{\text{\tiny 69.4 - 74.5}}$ 39.337.5 - ${}_{\text{\tiny 37.5 - 40.9}}$ 70.469.2 - ${}_{\text{\tiny 69.2 - 71.4}}$ 46.745.4 - ${}_{\text{\tiny 45.4 - 47.9}}$
Table S10: In-domain (ID) and out-of-domain (OOD) performance of Llama-3.1-8B-Instruct variants fine-tuned on m23k and further trained on reasoning-hard, knowledge-hard, or random-hard examples using SFT or RL. Models are evaluated separately on knowledge and reasoning subsets to assess generalization. While all models show performance drops in the OOD setting, those trained with RL—especially on reasoning-hard examples—exhibit stronger generalization and maintain higher accuracy across both knowledge and reasoning tasks.
Model Knowledge Reasoning Overall
ID OOD ID OOD ID OOD
Qwen w/ m23k SFT 62.260.8 - ${}_{\text{\tiny 60.8 - 63.3}}$ 45.744.1 - ${}_{\text{\tiny 44.1 - 47.6}}$ 61.558.8 - ${}_{\text{\tiny 58.8 - 64.0}}$ 32.731.0 - ${}_{\text{\tiny 31.0 - 34.4}}$ 62.160.9 - ${}_{\text{\tiny 60.9 - 63.2}}$ 39.638.4 - ${}_{\text{\tiny 38.4 - 40.8}}$
+ Random Hard SFT 62.661.4 - ${}_{\text{\tiny 61.4 - 63.9}}$ 48.446.8 - ${}_{\text{\tiny 46.8 - 50.1}}$ 62.760.0 - ${}_{\text{\tiny 60.0 - 65.5}}$ 34.532.9 - ${}_{\text{\tiny 32.9 - 36.2}}$ 62.661.4 - ${}_{\text{\tiny 61.4 - 63.8}}$ 41.940.6 - ${}_{\text{\tiny 40.6 - 43.0}}$
+ Knowledge Hard SFT 63.862.5 - ${}_{\text{\tiny 62.5 - 65.0}}$ 47.745.9 - ${}_{\text{\tiny 45.9 - 49.4}}$ 61.859.2 - ${}_{\text{\tiny 59.2 - 64.5}}$ 32.731.0 - ${}_{\text{\tiny 31.0 - 34.4}}$ 63.462.3 - ${}_{\text{\tiny 62.3 - 64.7}}$ 40.639.4 - ${}_{\text{\tiny 39.4 - 41.8}}$
+ Reason Hard SFT 64.262.8 - ${}_{\text{\tiny 62.8 - 65.5}}$ 48.146.3 - ${}_{\text{\tiny 46.3 - 49.8}}$ 65.362.5 - ${}_{\text{\tiny 62.5 - 68.2}}$ 36.034.3 - ${}_{\text{\tiny 34.3 - 37.8}}$ 64.463.2 - ${}_{\text{\tiny 63.2 - 65.6}}$ 42.441.2 - ${}_{\text{\tiny 41.2 - 43.6}}$
+ Random Hard RL 63.962.6 - ${}_{\text{\tiny 62.6 - 65.2}}$ 50.148.4 - ${}_{\text{\tiny 48.4 - 51.9}}$ 64.962.2 - ${}_{\text{\tiny 62.2 - 67.5}}$ 34.432.7 - ${}_{\text{\tiny 32.7 - 36.1}}$ 64.162.9 - ${}_{\text{\tiny 62.9 - 65.3}}$ 42.741.6 - ${}_{\text{\tiny 41.6 - 44.0}}$
+ Knowledge Hard RL 64.262.8 - ${}_{\text{\tiny 62.8 - 65.5}}$ 47.345.5 - ${}_{\text{\tiny 45.5 - 49.0}}$ 62.359.4 - ${}_{\text{\tiny 59.4 - 65.0}}$ 34.332.6 - ${}_{\text{\tiny 32.6 - 36.1}}$ 63.862.6 - ${}_{\text{\tiny 62.6 - 65.0}}$ 41.240.0 - ${}_{\text{\tiny 40.0 - 42.4}}$
+ Reason Hard RL 64.863.4 - ${}_{\text{\tiny 63.4 - 66.0}}$ 48.847.1 - ${}_{\text{\tiny 47.1 - 50.5}}$ 66.163.5 - ${}_{\text{\tiny 63.5 - 68.8}}$ 35.834.1 - ${}_{\text{\tiny 34.1 - 37.5}}$ 65.063.9 - ${}_{\text{\tiny 63.9 - 66.2}}$ 42.741.4 - ${}_{\text{\tiny 41.4 - 43.9}}$
Table S11: In-domain (ID) and out-of-domain (OOD) performance of Qwen3-7B-Instruct variants fine-tuned on m23k and further trained on reasoning-hard, knowledge-hard, or random-hard examples using SFT or RL. Evaluation is broken down by knowledge and reasoning subsets to assess generalization. Models trained on reasoning-hard examples—especially with SFT—show the most improvement on reasoning tasks, including in the OOD setting. RL training further enhances robustness, with the Reason Hard RL variant achieving the strongest overall OOD performance.
Model MQA MCQA PQA MMLU MB4 MD5 MDX HLE GPQA Lct. NEJM
                    Knowledge Accuracy
m1-7B-23K 74.3 61.0 73.0 66.4 85.7 71.4 16.0 13.9 52.8 56.9 57.4
HuatuoGPT-o1-8B 71.1 58.2 79.9 59.1 42.9 71.4 14.9 19.1 51.7 60.9 61.7
OpenBioLLM-8B 46.6 50.3 58.9 47.3 42.9 57.1 11.2 13.9 32.8 43.4 48.0
UltraMedical-8B 54.8 53.2 77.8 52.7 42.9 57.1 12.6 14.8 39.2 52.0 54.1
MedReason-8B 67.5 60.6 79.2 61.3 71.4 57.1 19.2 25.2 47.5 58.0 59.3
BioMed-R1-8B 78.9 66.5 77.6 70.4 71.4 71.4 19.7 9.6 55.7 62.3 64.4
medgemma-27b-text-it 83.4 70.8 74.0 75.4 71.4 71.4 24.7 14.8 55.5 65.8 69.6
m1-32B-1K 84.5 66.5 76.5 77.8 71.4 57.1 23.3 15.7 60.8 66.9 66.6
HuatuoGPT-o1-70B 86.8 75.0 81.1 77.7 57.1 71.4 27.2 13.0 67.7 70.1 72.3
OpenBioLLM-70B 75.0 66.9 78.6 69.8 71.4 85.7 18.3 15.7 54.4 61.9 61.7
UltraMedical-70B 80.2 71.7 80.0 75.4 71.4 71.4 22.0 10.5 59.2 72.6 69.3
BioMed-R1-32B 87.0 69.5 77.5 80.1 85.7 71.4 26.0 9.6 69.3 69.8 68.7
                    Reasoning Accuracy
m1-7B-23K 67.9 64.7 90.5 67.0 62.5 56.1 18.1 9.4 46.7 64.9 63.5
HuatuoGPT-o1-8B 71.2 63.6 90.5 72.1 53.8 53.2 16.6 6.2 33.3 61.1 66.4
OpenBioLLM-8B 47.1 49.3 66.7 45.3 34.9 31.9 12.3 18.8 33.3 50.4 51.5
UltraMedical-8B 58.1 56.3 85.7 59.2 53.8 48.8 14.2 12.5 40.0 58.8 59.5
MedReason-8B 63.3 60.2 90.5 76.0 55.5 52.5 21.2 12.5 0.0 61.8 63.1
BioMed-R1-8B 73.2 61.9 92.9 77.1 69.4 63.1 20.4 3.1 33.3 61.8 72.6
medgemma-27b-text-it 77.2 70.9 78.6 78.2 77.1 67.1 25.3 3.1 60.0 67.9 75.9
m1-32B-1K 77.8 68.9 88.1 78.8 70.8 68.1 25.2 18.8 13.3 74.0 75.2
HuatuoGPT-o1-70B 82.8 81.0 92.9 84.9 77.7 72.1 25.1 6.2 33.3 73.3 77.0
OpenBioLLM-70B 68.7 67.8 88.1 78.2 59.1 51.5 19.0 6.2 33.3 66.4 70.4
UltraMedical-70B 76.6 72.0 88.1 78.8 68.4 64.1 21.0 3.1 20.0 74.8 73.0
BioMed-R1-32B 83.0 72.8 90.5 84.9 74.1 70.8 26.4 3.1 33.3 72.5 79.2
                    Overall Accuracy
m1-7B-23K 70.1 61.3 73.7 66.5 63.0 56.5 17.4 12.9 52.6 59.5 60.2
HuatuoGPT-o1-8B 71.2 58.6 80.3 60.7 53.6 53.6 16.0 16.3 51.0 60.9 63.8
OpenBioLLM-8B 46.9 50.2 59.2 47.0 35.1 32.5 12.0 15.0 32.8 45.6 49.6
UltraMedical-8B 57.0 53.4 78.1 53.5 53.6 49.0 13.7 14.3 39.2 54.1 56.6
MedReason-8B 64.7 60.6 79.7 63.0 55.8 52.6 20.5 22.4 45.6 59.2 61.0
BioMed-R1-8B 75.2 66.1 78.2 71.1 69.5 63.3 20.2 8.2 54.9 62.1 68.2
medgemma-27b-text-it 79.3 70.8 74.2 75.8 76.9 67.2 25.1 12.2 55.6 66.5 72.5
m1-32B-1K 80.1 66.7 77.0 77.9 70.8 67.9 24.6 16.3 59.0 69.2 70.5
HuatuoGPT-o1-70B 84.2 75.5 81.6 78.5 77.3 72.1 25.8 11.6 66.4 71.1 74.5
OpenBioLLM-70B 70.9 67.0 79.0 70.8 59.4 52.3 18.8 13.6 53.6 63.3 65.7
UltraMedical-70B 77.8 71.7 80.3 75.7 68.5 64.3 21.3 8.9 57.7 73.3 71.0
BioMed-R1-32B 84.4 69.8 78.0 80.7 74.4 70.8 26.3 8.2 67.9 70.6 73.5
Table S12: Benchmark-wise performance of all evaluated models across twelve biomedical and general-domain reasoning datasets. We report knowledge, reasoning accuracy, and overall accuracy for each model. Our proposed models—BioMed-R1-8B and BioMed-R1-32B—demonstrate consistently strong performance across benchmarks, outperforming other open-source models of similar or larger scale.